-4(5p + 3) = -12 - 6p
Exactly how do I do this ?

Answer by lenny460(1073)

-4(5p + 3) = -12 - 6p
Exactly how do I do this ?
-4 * 5p = -20p
-4 * 3 = -12
-20p - 12 = -12 - 6p
Add 12 to each side of the equation.
-20p - 12 + 12 = -12 - 6p + 12
-20p = -6p
Add 6p to each side of the equation.
-20p + 6p = -6p + 6p
-14p = 0